ENR CTIO Blanco 4-m 2012-12-29

Date: 
Saturday, December 29, 2012
Telescope: 
CTIO Blanco 4-m
Submitter: 
David James
Submitter Email: 
djj [at] ctio [dot] noao [dot] edu
Observer Support: 
Hernan Tirado
Night Assistant: 
Claudio Aguilera
Also Present: 
Javier R.,
Problems: 

Several (>5) TCS failures tonight. 

Comments: 

Generally, pretty good night. 



Very bright Moon (couple of days past full) meant that we spent most of the night observing SNE fields. 



In fact, all SNe fields were observed except for X1, X2, C1 and C2. 



We targetted X3_i, E1, E2, S1, S2, C3_gri. 



At the end of the night, OBSTAC was running the show, with several tens of SPT-East fields observed. 



Standard star fields were observed at the beginning and end of the night. Early evening nautical-twilight standard fields were observed in a backwards filter order (in the wrong order !) to test the effects of decreasing sky-brightness as a function of photometric solution. 



We experienced several (at least 5 I think) TCS failures, with several occuring at one OBSTAC position (see elog entries #40280 and #40276 for observing times). 



There was a mountain-top power failure at about 00:30 local (elog entry #40224), and while the UPS kept us up and running, it crashed the TCS and the guider. Most admirably however, SISPI and DECam kept on observing right through it, even though the guider and TCS crashed. Obviously, there are a few strange images at this time because the telescope was not tracking (or guiding !). 



Operational mode for the coming night (30-DEC-2012) will be to observe the two remaining SNe fields that most desperately need epochs (X1 and X2), and then let OBSTAC do its thing ! 



DJJ Generally, pretty good night. 



Very bright Moon (couple of days past full) meant that we spent most of the night observing SNE fields. 



In fact, all SNe fields were observed except for X1, X2, C1 and C2. 



We targetted X3_i, E1, E2, S1, S2, C3_gri. 



At the end of the night, OBSTAC was running the show, with several tens of SPT-East fields observed. 



Standard star fields were observed at the beginning and end of the night. Early evening nautical-twilight standard fields were observed in a backwards filter order (in the wrong order !) to test the effects of decreasing sky-brightness as a function of photometric solution. 



We experienced several (at least 5 I think) TCS failures, with several occuring at one OBSTAC position (see elog entries #40280 and #40276 for observing times). 



There was a mountain-top power failure at about 00:30 local (elog entry #40224), and while the UPS kept us up and running, it crashed the TCS and the guider. Most admirably however, SISPI and DECam kept on observing right through it, even though the guider and TCS crashed. Obviously, there are a few strange images at this time because the telescope was not tracking (or guiding !). 



Operational mode for the coming night (30-DEC-2012) will be to observe the two remaining SNe fields that most desperately need epochs (X1 and X2), and then let OBSTAC do its thing ! 



DJJ Generally, pretty good night. 



Very bright Moon (couple of days past full) meant that we spent most of the night observing SNE fields. 



In fact, all SNe fields were observed except for X1, X2, C1 and C2. 



We targetted X3_i, E1, E2, S1, S2, C3_gri. 



At the end of the night, OBSTAC was running the show, with several tens of SPT-East fields observed. 



Standard star fields were observed at the beginning and end of the night. Early evening nautical-twilight standard fields were observed in a backwards filter order (in the wrong order !) to test the effects of decreasing sky-brightness as a function of photometric solution. 



We experienced several (at least 5 I think) TCS failures, with several occuring at one OBSTAC position (see elog entries #40280 and #40276 for observing times). 



There was a mountain-top power failure at about 00:30 local (elog entry #40224), and while the UPS kept us up and running, it crashed the TCS and the guider. Most admirably however, SISPI and DECam kept on observing right through it, even though the guider and TCS crashed. Obviously, there are a few strange images at this time because the telescope was not tracking (or guiding !). 



Operational mode for the coming night (30-DEC-2012) will be to observe the two remaining SNe fields that most desperately need epochs (X1 and X2), and then let OBSTAC do its thing ! 



DJJ 

Conditions: 
General Conditions
Temperature Max: 
18.00
Temperature Min: 
15.00
Humidity Max: 
40
Humidity Min: 
5
Wind Speed Max: 
6.00
Wind Speed Min: 
1.00
Comment: 
Very good night -- 8-oktas clear

Quarters

Program: 
Proposal: 
2012B-0003
Target of Oportunity: 
No
Time Observed: 
7.0
Time Engineering: 
0.0
Time lost technical: 
1.0
Time lost weather: 
0.0
Time lost other: 
0.0
Description: 
DES exSV
Primary Investigator: 
Josh Frieman
Institution: 
Fermilab
Instrument: 
CTIO 4m DECam
Organization: 
N/A
Total Program Time: 
8
Total Times Time Observed Time Engineering Time lost technical Time lost weather Time lost other Total Program Time
2012-12-29 7.0 0.0 1.0 0.0 0.0 8